# Computational Advantage in Hybrid Quantum Neural Networks: Myth or Reality?

### v2

@@ I Introduction
Quantum Machine Learning (QML) is an emerging research area that integrates quantum computing with machine learning (ML) techniques, aiming to enhance computational power and efficiency in processing complex datasets [1, 2, 3, 4]. Quantum Neural Networks (QNNs) are a core component of QML, employing quantum circuits/layers to mimic the structure and function of classical neural networks (NNs)[5, 6]. Hybrid Quantum Neural Networks (HQNNs) advance this concept by combining quantum and classical NN, leveraging the strengths of both quantum and classical computation[7, 8, 9]. HQNNs are particularly advantageous in the Noisy Intermediate-Scale Quantum (NISQ) era, where they can capitalize on quantum capabilities for specific sub-tasks while relying on classical resources to manage error-prone operations [10, 11]. In a typical HQNN architecture, one or more hidden layers of a classical NN are replaced by a trainable quantum layer [12, 13], as shown in Fig. 1.
@@ I-B Our Contributions
Summary of key results: This study demonstrates the superior scalability of HQNNs compared to classical NNs as problem complexity increases. In our analysis, we find that despite simulation overhead, HQNNs with expressive quantum layers require approximately $7.5\%$ fewer FLOPs and $42.3\%$ fewer parameters to solve complex problems ($110$ features in our case) than classical networks. Additionally, HQNNs scale more efficiently, with a $53.1\%$ increase in FLOPs while going from simple to complex tasks, compared to $88.5\%$ for classical models. Similarly, the rate of increase in number of parameters for classical models scales is $88.5\%$, while HQNNs show a slower rate of increase in FLOPs that accounts to $81.4\%$.
@@ II Complexity of Neural Networks
There is no single, definitive measure for assessing the NN’s complexity, however, several approximate metrics exist[31, 32, 33, 34, 35, 36]. Therefore, we argue that relying on a single measure may not provide a comprehensive understanding of the network’s complexity. Hence, it is advisable to use multiple methods to better estimate the complexity. We use FLOPs and the number of parameters as indicators to assess the network’s complexity.
@@ III-B Exhaustive Classical Model Search
Classical Model Search Space: We restrict the classical models to have a maximum of $n=3$ layers, with the number of neurons in each layer chosen from the set $m=\{2,4,6,8,10\}$ resulting in a search space of total $155$ model combinations for classical models for each complexity level.
@@ III-C Exhaustive Hybrid Model Search
For hybrid models, the number of neurons in the first and last classical layers are fixed. The number of neurons in the input layer are equal to the number of qubits because we use angle encoding that requires one qubit for one feature[45]. The neurons in the output layer are equal to the number of classes ($3$) in the dataset.
@@
Quantum Layer Design: For the underlying quantum layers in HQNNs, we use two different types designs which are are separately evaluated: the Strongly Entangling Layer (SEL) and the Basic Entangling Layer (BEL), as shown in Fig. 5. For both layer types, each qubit number is experimented with all the layer depths, to identify the optimal configuration, i.e., that meets the preset accuracy threshold of $\geq 90\%$.
@@ III-F Models Training
The sorted list of models are trained sequentially for various levels of problem complexity. To account for randomness of parameter initialization and statistical reliability of the results, each model combination subjected to five independent runs. Each run involves training for $100$ epochs, and the highest training and validation accuracy across epochs is recorded. The maximum accuracies from all runs are averaged to provide a robust performance measure for each model combination for a given complexity level. If any model achieves an average training and validation accuracy of $\geq 90\%$, further evaluation of additional model combinations for that complexity level is stopped. This entire process is repeated five times to address the stochastic nature of NNs training, where different model configurations might result as optimal. By repeating the procedure, a more accurate estimate of the average complexity required to solve the problem is obtained. Once five top-performing combinations are identified for a given complexity level, the evaluation continues with the next complexity level.
@@ IV-A Classical Models: Analysis of FLOPs Consumption
In this section, we present our findings on the computational complexity required by classical models to solve the problem under predefined accuracy conditions, as shown in Fig. 6. By varying problem complexity (increasing features and noise, see Section III-A), we evaluate the scalability and computational efficiency of classical models. We observe that at lower complexity, simpler models achieve the desired accuracy. However, as the problem complexity increases (e.g., up to $110$ and associated high noise level), more sophisticated models, i.e., with additional layers and neurons are needed, resulting in a significant increase in FLOPs and hardware resources.
@@ V Conclusion
(A2) Our analysis reveals that as the problem complexity increases, classical NNs demonstrate a substantial rise in both FLOPs and the number of parameters due to the requirement of more sophisticated architecture to maintain a certain accuracy. This indicates limited adaptability of classical NNs to the increasing complexity of the problem. In contrast, HQNNs exhibit a more efficient computational footprint. Specifically, quantum layers within HQNNs consume significantly fewer FLOPs and require fewer parameters as problem complexity grows, compared to classical models. Additionally, HQNNs adapt well to the growing problem complexity since the rate of increase in FLOPs and parameter count (while going from low to high complexity of the problem) is significantly lower in HQNNs compared to classical models. These findings suggest that HQNNs are better suited to adapting to complex problem scenarios, highlighting their potential advantages over classical models in managing computational complexity.

### v3

@@ I Introduction
Quantum Machine Learning (QML) is an emerging research area that integrates quantum computing with machine learning (ML) techniques, aiming to enhance computational power and efficiency in processing complex datasets [1, 2, 3, 4]. Quantum Neural Networks (QNNs) are a core component of QML, employing quantum circuits/layers to mimic the structure and function of classical neural networks (NNs)[5, 6, 7]. Hybrid Quantum Neural Networks (HQNNs) advance this concept by combining quantum and classical NN, leveraging the strengths of both quantum and classical computation[8, 9, 10]. HQNNs are particularly advantageous in the Noisy Intermediate-Scale Quantum (NISQ) era, where they can capitalize on quantum capabilities for specific sub-tasks while relying on classical resources to manage error-prone operations [11, 12]. In a typical HQNN architecture, one or more hidden layers of a classical NN are replaced by a trainable quantum layer [13, 14], as shown in Fig. 1.
@@ I-B Our Contributions
Summary of key results: This study demonstrates the superior scalability of HQNNs compared to classical NNs as problem complexity increases. In our analysis, we find that despite simulation overhead, HQNNs with expressive quantum layers require approximately $7.5\%$ fewer FLOPs and $42.3\%$ fewer parameters to solve complex problems ($110$ features in our case) than classical networks. Additionally, HQNNs scale more efficiently, with a $53.1\%$ increase in FLOPs consumption while going from simple to complex tasks, compared to $88.5\%$ increase in FLOPs for classical models. Similarly, the rate of increase in number of parameters for classical models is $88.5\%$, while HQNNs show a slower rate of increase in parameters that accounts to $81.4\%$.
@@ II Complexity of Neural Networks
There is no single, definitive measure for assessing the NN’s complexity, however, several approximate metrics exist[32, 33, 34, 35, 36, 37]. Therefore, we argue that relying on a single measure may not provide a comprehensive understanding of the network’s complexity. Hence, it is advisable to use multiple methods to better estimate the network’s complexity for a given task. We use FLOPs and the number of parameters as indicators to assess the network’s complexity.
@@ III-B Exhaustive Classical Model Search
Classical Model Search Space: We restrict the classical models to have a maximum of $n=3$ layers, with the number of neurons in each layer chosen from the set $m=\{2,4,6,8,10\}$ ($m=5$), resulting in a search space of total $155$ model combinations for classical models for each complexity level.
@@ III-C Exhaustive Hybrid Model Search
For hybrid models, the number of neurons in the first and last classical layers are fixed. The number of neurons in the input layer are equal to the number of qubits because we use angle encoding that requires one qubit to encode one feature[46]. The neurons in the output layer are equal to the number of classes ($3$) in the dataset.
@@
Quantum Layer Design: For the underlying quantum layers in HQNNs, we use two different types designs which are separately evaluated: the Strongly Entangling Layer (SEL) and the Basic Entangling Layer (BEL), as shown in Fig. 5. For both layer types, each qubit number is experimented with all the layer depths, to identify the optimal configuration, i.e., that meets the preset accuracy threshold of $\geq 90\%$.
@@ III-F Models Training
The sorted list of models are trained sequentially for various levels of problem complexity. To account for the inherent randomness in parameter initialization and to ensure the statistical reliability of the results, each model configuration is evaluated over five independent runs. This approach mitigates the effects of stochastic variations and provides a more robust assessment of model performance across different initialization instances. Each run involves training for $100$ epochs, and the highest training and validation accuracy across epochs is recorded. The maximum accuracies from all runs are averaged to provide a robust performance measure for each model combination for a given complexity level. If any model achieves an average training and validation accuracy of $\geq 90\%$, further evaluation of additional model combinations for that complexity level is stopped. This entire process is repeated five times to address the stochastic nature of NNs training, where different model configurations might result as optimal at different time instances. By repeating the procedure, a more accurate estimate of the average complexity required to solve the problem is obtained. Once five top-performing combinations are identified for a given complexity level, the evaluation continues with the next complexity level.
@@ IV-A Classical Models: Analysis of FLOPs Consumption
In this section, we present our findings on the computational complexity (in terms of FLOPs consumption) required by classical models to solve the problem under predefined accuracy conditions, as shown in Fig. 6. By varying the problem complexity (increasing features and noise, see Section III-A), we evaluate the scalability and computational efficiency of classical models. We observe that at lower complexity, simpler models achieve the desired accuracy. However, as the problem complexity increases (e.g., up to $110$ and associated high noise level), more sophisticated models, i.e., with additional layers and neurons are needed, resulting in a significant increase in FLOPs and hardware resources.
@@ V Conclusion
(A2) Our analysis reveals that as the problem complexity increases, classical NNs demonstrate a substantial rise in both FLOPs and the number of parameters due to the requirement of more sophisticated architecture to maintain a certain accuracy. This indicates limited adaptability of classical NNs to the increasing complexity of the problem. In contrast, HQNNs exhibit a more efficient computational footprint. Specifically, quantum layers within HQNNs consume significantly fewer FLOPs and require fewer parameters as problem complexity grows, compared to classical models. Additionally, HQNNs adapt well to the growing problem complexity since the rate of increase in FLOPs and parameter count (while going from low to high complexity of the problem) is significantly lower in HQNNs compared to classical models. These findings indicate that HQNNs exhibit superior adaptability to complex problem scenarios, highlighting their potential advantages over classical models in managing computational complexity.
@@
Acknowledgements
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008.